OpenStudy (anonymous):

Which expression is equivalent to (53)−2? 1 over 5 to the 3rd negative 1 over 5 times 5 times 5 times 5 times 5 times 5 times 53 1 over 5 times 5 times 5 times 5 times 5 times 5 times

OpenStudy (anonymous):

Do you mean \[53^{-2}\]

OpenStudy (anonymous):

its 5 with the power of 3

OpenStudy (anonymous):

Anything raised to a negatve exponent is equivalent to it being a fraction like so. These are a few examples \[3^{-3} = 1/3^{3} \rightarrow 2^{-15}=1/2^{15}\]

OpenStudy (anonymous):